kmyahhayes1234:

Determine the x-intercept of the line whose equation is given: y = StartFraction x Over 2 EndFraction minus 3 a. (6, 0) b. (negative 6, 0) c. (0, three-halves) d. (Negative three-halves, 0)

Angle:

Hello @kmyahhayes1234 and Welcome to QuestionCove! Give me a moment to read your question :)

Angle:

Ok, so I believe your equation looks like this: \(y = \frac{x}{2}-3\) What we are looking for is the "x-intercept" What this means is that "What does x equal when y is zero?" so, let's try to plug in y = 0 into our equation and figure out what x is \(0 = \frac{x}{2}-3\) how can we solve for x?

Angle:

@kmyahhayes1234 if you are confused by anything I have just said, just let me know! I want to help :)

jhonyy9:

so i think use formula \[y = mx +b \] \[y = \frac{ 1 }{ 2 }x -3\]

jhonyy9:

what @Angle 've wrote above to you too